# Measurement as Sheafification: Context, Logic, and Truth after Quantum Mechanics

### v1

@@ Abstract
Quantum measurement is commonly posed as a dynamical tension between linear Schrödinger evolution and an ad hoc collapse rule. I argue that the deeper conflict is logical: quantum theory is inherently contextual, whereas the classical tradition presupposes a single global, Boolean valuation. Building on Bohr’s complementarity, the Einstein–Podolsky–Rosen argument and Bell’s theorem, I recast locality and completeness as the existence of a global section of a presheaf of value assignments over the category of measurement contexts. The absence of global sections expresses the impossibility of context-independent description, and Čech cohomology measures the resulting obstruction. The internal logic of the presheaf topos is intuitionistic, and the seven-valued contextual logic proposed by Ghose and Patra is exhibited as a finite Heyting algebra capturing patterns of truth, falsity and indeterminacy across incompatible contexts. Classical physics corresponds to the sheaf case, where compatible local data glue and Boolean logic is effectively restored. Measurement is therefore reinterpreted as sheafification of presheaf-valued truth rather than as a physical breakdown of unitarity. Finally, a $\sigma$$\lambda$ dynamics motivated by stochastic mechanics provides a continuous interpolation between strongly contextual and approximately classical regimes, dissolving the usual measurement paradoxes and apparent nonlocality as artefacts of an illegitimate demand for global truth.
@@
Keywords: quantum measurement; contextuality; presheaves and sheaves; topos/Heyting logic; $\sigma$$\lambda$ dynamics
1 Introduction: Why the Measurement Problem Refuses to Go Away
From its very inception, quantum mechanics has contained an unresolved conceptual tension at its core. In von Neumann’s canonical formulation [1], the theory requires two fundamentally different kinds of time evolution. On the one hand, there is the continuous, deterministic, and linear evolution governed by the Schrödinger equation. This evolution is fully quantum mechanical, unitary, and applies universally to all isolated systems. On the other hand, there is the discontinuous and stochastic “collapse of the wavefunction” associated with measurement, by which a superposition of possible outcomes is replaced by a single definite result. This second process is not described by the Schrödinger equation and is not governed by the same physical principles.
The tension between these two processes lies at the heart of what has come to be called the “measurement problem.” If the quantum formalism is taken to apply universally, then measurement devices themselves ought to evolve unitarily, leading only to ever larger superpositions rather than definite outcomes. If, however, one insists on definite outcomes, as one must on empirical grounds, an additional, non-unitary intervention seems unavoidable. The theory thus appears to require two incompatible modes of description, without specifying when or how the transition between them occurs.
The measurement problem has spawned many and varied interpretations, including the de Broglie–Bohm approach [2], the Everett many-worlds view [3], the decohering-histories programme of Omnès [4], Rovelli’s relational interpretation [5], QBism [6], and others. Yet the problem has persisted, not because of a lack of technical ingenuity or experimental guidance, but because, in my view, it is typically framed within an inherited classical conception of logic. From the earliest formulations of quantum mechanics, the formalism has insisted on context-dependence, complementarity, and the incompatibility of certain types of experimental arrangements, as repeatedly emphasized by Bohr [7]. Interpretation has nevertheless attempted to force these features back into a framework of globally valid truth and context-independent description. The resulting tension manifests as the so-called “collapse of the wavefunction,” understood as a physical discontinuity that must somehow be added to otherwise continuous dynamics.
The persistence of the problem becomes especially transparent in paradoxes such as Wigner’s friend [8]. In this scenario, one observer inside a laboratory performs a measurement and records a definite outcome, while a second observer, who treats the entire laboratory quantum mechanically, assigns a superposed state to the combined system of apparatus, observer, and measured object. Both descriptions are sanctioned by the formalism, yet they cannot be jointly embedded within a single, globally consistent account. The paradox does not lie in experimental ambiguity, but in the coexistence of incompatible yet equally legitimate descriptions.
Attempts to resolve this tension typically require the introduction of something external to the quantum formalism itself: a distinguished role for observers, a fundamental boundary between classical and quantum domains, or even an appeal to consciousness. Such moves underscore the central difficulty: the theory resists a single, context-independent notion of truth. Wigner’s friend brings into sharp focus the need to rethink not the dynamics of quantum mechanics, but the logical framework within which its physical statements are interpreted.
This paper advances the hypothesis that the measurement problem is not fundamentally a problem of physics, but of logic: it arises from insisting that a single, Boolean logical structure should govern both quantum and classical domains. Once logic itself is allowed to be conditioned by physical context, the problem changes character. What appears as a mysterious physical collapse may instead be understood as a transition between distinct logical regimes. The central aim here is to articulate this transition in conceptual terms, using the modern language of category theory (contexts, presheaves, and sheaves), and to show how quantum mechanics itself motivates such a rethinking.
@@ 8 Dynamics as a Continuous Passage Between Logical Regimes
Nelson begins from the classical Newtonian picture, but assumes that particles undergo a Brownian motion in a universal medium with a diffusion constant $\sigma$ superposed on their regular motion. Instead of a single velocity field, one has forward and backward mean velocities, and the particle trajectories are described by stochastic differential equations. The key result is that, under suitable assumptions, the ensemble dynamics of such stochastic trajectories can be recast into the familiar Schrödinger equation. The Planck constant $\hbar$ then appears not as a primitive quantum postulate, but as a parameter fixed by the relation
@@
where $m$ is the mass of the particle and $\sigma$ is the diffusion parameter characterising the underlying Brownian motion. [^1] More standardly one writes $\nu=\hbar/(2m)$ for the diffusion constant; the present notation absorbs numerical factors into $\sigma$ for simplicity.
@@
In the stochastic picture, the parameter $\lambda$ can be regarded as a function of the underlying diffusion, $\lambda=\lambda(\sigma)$, encoding how strongly the quantum potential survives coarse-graining and environmental influence. Large diffusion (or suitable choices of $m$ and $\sigma$ satisfying $\hbar=m\sigma$) support $\lambda\approx 1$ and hence fully quantum behaviour; as the effective diffusion weakens, the quantum potential term is diminished and $\lambda$ tends toward 0. The $\sigma$$\lambda$ dynamics thus implements a continuous deformation of the underlying Hamilton–Jacobi structure between quantum and classical limits. Davidson has argued that variational formulations of stochastic mechanics allow non-unique (even variable) diffusion, which dovetails with treating $\sigma$ as a tunable “quantumness” parameter [28].
@@ 11 Summary and Outlook
The analysis developed in this paper has suggested that the measurement problem in quantum mechanics is not primarily a dynamical paradox, but a manifestation of a deeper logical tension. The standard formulation tacitly presupposes that physical reality must be describable by a single, context-independent, globally valid Boolean valuation. Quantum theory, by contrast, forces upon us a world in which experimental arrangements are mutually incompatible and in which no global assignment of sharp values to all observables is possible. The apparent conflict between unitary evolution and wavefunction collapse, and the familiar worries about nonlocal “spooky action-at-a-distance”, are symptoms of this mismatch between a classical logical ideal and the contextual structure of the quantum formalism itself.
By making measurement contexts explicit and organising them into a category, we have seen that presheaves provide a natural semantics for quantum truth. Context-dependent value assignments form presheaves on the category of contexts; the failure of a global section becomes a precise expression of the breakdown of classical, context-independent description. The internal logic of the resulting presheaf topos is intuitionistic rather than Boolean, and concrete schemes such as the seven-valued logic proposed by Ghose and Patra can be understood as finite Heyting algebras encoding distinct patterns of truth, falsity and indeterminacy across incompatible contexts. Sheaves then mark the regime in which local data can be glued into global descriptions: classical physics appears as the special case in which presheaf data satisfy a sheaf condition and Boolean logic is effectively restored.
@@
Several directions for further work suggest themselves. On the conceptual side, one may seek a more systematic classification of contextual logics within presheaf topoi and a clearer comparison with other approaches to quantum logic, including the Birkhoff–von Neumann programme and various topos-theoretic reconstructions. On the mathematical side, it would be natural to refine the cohomological analysis of contextuality and to explore more fully how the $\sigma$$\lambda$ dynamics controls the transition between nontrivial and trivial cohomology. On the physical side, one may study concrete models—for instance, simple interferometric or spin systems—in which the deformation from presheaf-like to sheaf-like behaviour can be analysed quantitatively, and investigate whether intermediate regimes admit experimental signatures. Empirically, micellar aqueous surfactant systems have been reported to exhibit oscillatory behaviour between quantum-like and classical-like regimes, and can be modelled using an interpolating Schrödinger dynamics with continuously varying $\lambda$, providing suggestive evidence for a smooth quantum–classical passage [30]. Whatever the outcome of such developments, the central lesson remains: logic is not a neutral backdrop for physics, but an emergent structure reflecting the way our physical world organises its contexts.
@@ 12 Appendix A: Introduction to Category Theory
  • nonexistence subsubsection(of global truth-values.
@@ Acknowledgements
The author acknowledges the use of OpenAI’s ChatGPT for language polishing and assistance with LaTeX formatting and manuscript preparation. All scientific arguments, interpretations, and remaining errors are the author’s responsibility.

### v2

@@ Abstract
Quantum measurement is commonly posed as a dynamical tension between linear Schrödinger evolution and an ad hoc collapse rule. I argue that the deeper conflict is logical: quantum theory is inherently contextual, whereas the classical tradition presupposes a single global, Boolean valuation. Building on Bohr’s complementarity, the Einstein–Podolsky–Rosen argument and Bell’s theorem, I recast locality and completeness as the existence of a global section of a presheaf of value assignments over the category of measurement contexts. The absence of global sections expresses the impossibility of context-independent description, and Čech cohomology measures the resulting obstruction. The internal logic of the presheaf topos is intuitionistic, and the Ghose–Patra seven-valued contextual logic is exhibited as a finite Heyting algebra capturing patterns of truth, falsity and indeterminacy across incompatible contexts. Classical physics corresponds to the sheaf case, where compatible local data glue and Boolean logic is effectively restored. Measurement is therefore reinterpreted as sheafification of presheaf-valued truth rather than as a physical breakdown of unitarity. Finally, a $\sigma$$\lambda$ dynamics motivated by stochastic mechanics provides a continuous interpolation between strongly contextual and approximately classical regimes, dissolving the usual measurement paradoxes and apparent nonlocality as artefacts of an illegitimate demand for global truth.
@@
Keywords: quantum measurement; contextuality; presheaves and sheaves; topos/Heyting logic; Saptabhaṅgī; $\sigma$$\lambda$ dynamics
@@ 7 Dynamics as a Continuous Passage Between Logical Regimes
Nelson begins from the classical Newtonian picture, but assumes that particles undergo a Brownian motion with a diffusion constant $\sigma$ superposed on their regular motion. Instead of a single velocity field, one has forward and backward mean velocities, and the particle trajectories are described by stochastic differential equations. The key result is that, under suitable assumptions, the ensemble dynamics of such stochastic trajectories can be recast into the familiar Schrödinger equation. The Planck constant $\hbar$ then appears not as a primitive quantum postulate, but as a parameter fixed by the relation
@@
where $m$ is the mass and $\sigma$ is the diffusion parameter characterising the underlying Brownian motion. [^1] More standardly one writes $\nu=\hbar/(2m)$ for the diffusion constant; the present notation absorbs numerical factors into $\sigma$ for simplicity.
@@
In the stochastic picture, the parameter $\lambda$ can be regarded as a function of the underlying diffusion, $\lambda=\lambda(\sigma)$, encoding how strongly the quantum potential survives coarse-graining and environmental influence. Large diffusion (or suitable choices of $m$ and $\sigma$ satisfying $\hbar=m\sigma$) support $\lambda\approx 1$ and hence fully quantum behaviour; as the effective diffusion weakens, the quantum potential term is diminished and $\lambda$ tends toward 0. The $\sigma$$\lambda$ dynamics thus implements a continuous deformation of the underlying Hamilton–Jacobi structure between quantum and classical limits.
@@ 10 Summary and Outlook
The analysis developed in this paper has suggested that the measurement problem in quantum mechanics is not primarily a dynamical paradox, but a manifestation of a deeper logical tension. The standard formulation tacitly presupposes that physical reality must be describable by a single, context-independent, globally valid Boolean valuation. Quantum theory, by contrast, forces upon us a world in which experimental arrangements are mutually incompatible and in which no global assignment of sharp values to all observables is possible. The apparent conflict between unitary evolution and wavefunction collapse, and the familiar worries about nonlocal “spooky action-at-a-distance”, are symptoms of this mismatch between a classical logical ideal and the contextual structure imposed by the quantum formalism itself.
By making measurement contexts explicit and organising them into a category, we have seen that presheaves provide a natural semantics for quantum truth. Context-dependent value assignments form presheaves on the category of contexts; the failure of a global section becomes a precise expression of the breakdown of classical, context-independent description. The internal logic of the resulting presheaf topos is intuitionistic rather than Boolean, and concrete schemes such as the Ghose–Patra seven-valued logic can be understood as finite Heyting algebras encoding distinct patterns of truth, falsity and indeterminacy across incompatible contexts. Sheaves then mark the regime in which local data can be glued into global descriptions: classical physics appears as the special case in which presheaf data satisfy a sheaf condition and Boolean logic is effectively restored.
@@
Quantum gravity and the sheafification of spacetime.
A natural question is how the present perspective bears on quantum gravity. In the canonical approach, the central condition is the Wheeler–DeWitt constraint, $\widehat{\mathcal{H}}\Psi=0$, understood as the quantization of the Hamiltonian constraint of general relativity [29]. Although this is a manifestly quantum equation, its semiclassical (Born–Oppenheimer/WKB) expansion provides a standard bridge to classical spacetime: for a wavefunctional of the form $\Psi[h]\sim A[h]\exp(iS[h]/\hbar)$, the leading order yields the Einstein–Hamilton–Jacobi equation for $S[h]$, reproducing classical Einstein dynamics for a chosen geometric branch, with higher orders describing quantum fields on that background [30, 31]. From the present viewpoint, the recovery of classical general relativity may therefore be read as a two-step emergence: first, restriction to an appropriate quasi-classical branch within the constraint-satisfying sector; second, restoration of descent so that locally defined geometric data (chart, frame or connection descriptions on overlapping regions) glue into a globally communicable spacetime geometry. This use of gluing as a diagnostic of “classicality” resonates with earlier proposals to bring topos and sheaf ideas to quantum theory and quantum gravity [32] and with finitary sheaf approaches aimed at approximating continuum spacetime structure [33]. In our setting, the distinctive claim is that such restoration of global gluing can be tied to a continuous classicalization controlled by the $\sigma$$\lambda$ dynamics.
Several directions for further work suggest themselves. On the conceptual side, one may seek a more systematic classification of contextual logics within presheaf topoi and a clearer comparison with other approaches to quantum logic, including the Birkhoff–von Neumann programme and various topos-theoretic reconstructions. On the mathematical side, it would be natural to refine the cohomological analysis of contextuality and to explore more fully how the $\sigma$$\lambda$ dynamics controls the transition between nontrivial and trivial cohomology. On the physical side, one may study concrete models—for instance, simple interferometric or spin systems—in which the deformation from presheaf-like to sheaf-like behaviour can be analysed quantitatively, and investigate whether intermediate regimes admit experimental signatures. Whatever the outcome of such developments, the central lesson remains: logic is not a neutral backdrop for physics, but an emergent structure reflecting the way our physical world organises its contexts.
@@ 11 Appendix A: Introduction to Category Theory
A Čech-cohomological Obstruction to Contextuality
Let $\{C_{i}\to C\}_{i\in I}$ be a family of (overlapping) contexts covering a given context $C$ (e.g. the maximal compatible measurement contexts of an experimental scenario). Write $C_{ij}:=C_{i}\cap C_{j}$ and $C_{ijk}:=C_{i}\cap C_{j}\cap C_{k}$ for pairwise and triple overlaps.
To speak of Čech cohomology one works with an abelian coefficient presheaf. A convenient choice—used explicitly in cohomological contextuality—is to pass from a Set-valued presheaf $F$ to the presheaf of free abelian groups $\mathbb{Z}[F]$ generated by its sections, so that formal differences of restricted sections are meaningful [34].
A choice of local data is a Čech 0-cochain
$s=(s_{i})_{i\in I},\qquad s_{i}\in F(C_{i}).$
Its coboundary is the $1$-cochain $(\delta s)=(\delta s)_{ij}$ given on overlaps by
$(\delta s)_{ij}\;:=\;s_{j}|_{C_{ij}}\;-\;s_{i}|_{C_{ij}}\;\in\;\mathbb{Z}[F](C_{ij}).$
The sheaf gluing condition is precisely $\delta s=0$ (pairwise agreement on overlaps), in which case the family $\{s_{i}\}$ pastes to a global section $s\in F(C)$.
When $\delta s\neq 0$, the local sections fail to glue. Since $\delta^{1}\!\circ\delta^{0}=0$, the cochain $\delta s$ is automatically a $1$-cocycle, and its cohomology class
$\delta s]\;\in\;\check{H}^{1}(\{C_{i}\},\mathbb{Z}[F])$
is the obstruction class: if $\delta s]\neq 0$ then there is no global section compatible with the given local data.
In the Abramsky–Brandenburger sheaf-theoretic framework, an empirical model is a compatible family of probability distributions on maximal contexts, and contextuality is the non-existence of a global section explaining these marginals [18]. Abramsky, Mansfield and Barbosa define a Čech-cohomological obstruction class (built from an abelian presheaf derived from the support of the model) which vanishes whenever a global section exists; hence non-vanishing provides a robust sufficient witness of contextuality [34, 35].
@@
  • nonexistence of global truth-values.
@@ Acknowledgements
The author acknowledges the use of ChatGPT (OpenAI) for language polishing and assistance with formatting. The author takes full responsibility for the scientific content, interpretations, and any remaining errors.